# Localized Distributional Robustness in Submodular Multi-Task Subset Selection

### v2

@@ 6.2 Saturate with Preference for Satellite Selection
We now turn our attention to the assessment of Algorithm 1, Saturate with Preference. The general setting of the experiments remains similar to the previous subsection, with the same constellation parameters of $175^{\circ}:240/12/$ 1, and the same objective functions $f^{1},\ldots,f^{6}$. We compare the performance of the selection made by Saturate with Preference, with that of the unmodified Submodular Saturation Algorithm, by looking at the values achieved by the two objective functions that are assigned the highest weight by the random sampling of $Q$. In essence, this allows us to evaluate whether adding the element of preference to SSA works as intended. Indeed, Figure 5 demonstrates that for the two objective functions with the highest priority, Saturate with Preference leads to the selection of a subset that consistently achieves a higher utility score in comparison to SSA. Indeed, Figure 5 demonstrates that Saturate with Preference consistently outperforms SSA in terms of the performance on the objective functions with the highest assigned priority.
Fig. 5: The average performances over fifteen runs of Saturate with Preference in comparison to the Submodular Saturation Algorithm on the two objective functions with the highest assigned weight. The highlighted areas indicate one tenth standard deviation. The results have been put through a moving average filter with window size 6.
@@ 6.3 Application to Online Submodular Optimization
In this subsection, we will assess the performance of the novel formulation (66), which we will call the time-robust (TR) formulation for the sake of brevity, as detailed in Section 5. In summary, the TR formulation aims, using a combination of a momentum-like weighing scheme of the time-varying objective functions in an online setting, along with the idea of relative entropy regularization, to reuse the same selections made in one step over multiple time steps. In this way, it aims to be more cost-efficient in settings where the selection of more diverse elements incurs additional costs.
As a baseline, we choose the standard approach of treating each individual objective function $f_{t}$ observed at time step $t$ as a separate problem and solve it using the Stochastic Greedy algorithm, without any consideration for the conservation of solutions over time steps. Figure 6 demonstrates the comparison of our proposed approach with the time-window size $t_{w}=5$ against the standard approach, in terms of the utility achieved, the wall-clock time taken, and the total number of distinct elements used, i.e., $\lvert\bigcup_{t=0}^{T}S_{t}\rvert$, where $S_{t}$ represents the solution constructed at time step $t$ of the simulation. The results indicate that TR achieves a comparable yet slightly lower utility in comparison to the standard approach. The wall-clock time taken fluctuates in the TR formulation, since with a time-window size of $5$, the algorithm only observes during four of every five time steps and plays its solution, performing all function evaluations on every five time steps, although the total wall-clock time taken is in the same range as the standard approach. However, the total number of distinct elements chosen by the TR formulation is much lower, using less than half the number of distinct elements in comparison to the standard approach.
Fig. 6: The average performances over fifteen runs of the two algorithms in the online task detailed in Section 6, 6.3. The comparison criteria are the utility achieved, the wall-clock time taken, and the total number of distinct elements used. Regular represents the standard approach of solving each objective function $f_{t}$ observed at time step $t$ independently of the previous objective functions observed, using the Stochastic Greedy algorithm. TR represents the time-robust formulation detailed in Section 5, with a time-window size $t_{w}=5$. The highlighted areas indicate one standard deviation. The results have been put through a moving average filter with size 6.
@@ 6.4 Practical Application: Image Summarization
As a final demonstration using a more day-to-day application involving machine learning and signal processing, we apply the proposed method to image summarization. We tackle the case detailed in [31], which, in short, involves selecting the most representative $K\leq\lvert N\rvert$ images out of a dataset $N=\{1,2,\ldots,\lvert N\rvert\}$ of indexed images. More formally, for a selection $S\subseteq N$ of images, one can evaluate the utility $f^{i}(S)$ of each image $i\in N$ through its similarity to the closest image in the selection by:
$f^{i}(S)=1-\min_{e\in S}d(i,e),$ (69)
where $d(i,e)$ designates some measure of distance between images $i$ and $e$. Replicating the setting in the aforementioned work [31], we use the $\lvert N\rvert=819$-image Pokemon dataset[55], using image embeddings calculated with an AlexNet[30] trained on the ImageNet dataset[56]. We use the normalized cosine distance
$d(i,e)=\frac{v_{i}\cdot v_{e}}{\lVert v_{i}\rVert\lVert v_{e}\rVert}+\min_{(e,e^{{}^{\prime}})\in N^{2}}\frac{v_{e}\cdot v_{e^{{}^{\prime}}}}{\lVert v_{e}\rVert\lVert v_{e^{{}^{\prime}}}\rVert}.$ (70)
The average results obtained from fifteen runs of the algorithms, where in each the reference distribution $Q$ is the uniform distribution, are demonstrated in Figure 7. We observe that our proposed Local algorithm outperforms Saturate for nearly all values of the cardinality constraint $K$ on the reference distribution and on the local worst-case distribution, while also taking significantly less computation time, as evidenced by the results of the previous experiments. Regarding the performance on the worst-case task, Saturate only outperforms our algorithm within the cardinality constraint range of $K=1$ to $K=5$, but then both algorithms show virtually the same performance, again, with the Local algorithm being much less computationally expensive.
Fig. 7: The average performances over fifteen runs of the two algorithms focused on optimizing the global worst-case task, and the local worst-case distribution as guided by the reference distribution, evaluated on the three criteria of reference distribution performance, worst-case task performance, and local worst-case distribution performance in the image summarization task of Section 6, Subsection 6.4. Local represents the relative entropy-regularized Stochastic Greedy algorithm solving our novel formulation, aiming for local worst-case distributional robustness in the neighborhood of the reference distribution. Saturate (Global) represents the Submodular Saturation Algorithm proposed in [19], aiming for global worst-case task robustness. The highlighted areas indicate one standard deviation. The results have been put through a moving average filter with window size 6.
@@ 7 Conclusion
We proposed the novel formulation of Problem (14) to find a solution that is locally distributionally robust in the neighborhood of a reference distribution, assign an importance score to each task in a multi-task objective, using various statistical distances as a regularizer. The $\ell^{1}$ and $\ell^{\infty}$ distances led to the proposal of the Saturate with Preference algorithm, which incorporates an element of preference into the standard Submodular Saturation Algorithm. Afterward, we demonstrated that using relative entropy as a regularizer, through duality, one can show that this novel formulation is equivalent to Problem (25). Then, we proved that this dual formulation gives rise to the composition of a monotone increasing function with another normalized, monotone nondecreasing submodular function, which can be optimized with standard methods such as the Stochastic Greedy algorithm, enjoying theoretical guarantees. We proposed an application of the relative entropy regularized to online submodular optimization, through the use of a momentum-like weighing of the objective functions observed at each time step. We then experimentally corroborated our results for all three of the proposed settings, motivated by a practical scenario involving a sensor selection problem within a simulation of LEO satellites, using weak-submodular functions. For the relative entropy regularization setting, we compared our algorithm with two other algorithms focused on optimizing the performance of the worst-case task, and on directly optimizing the performance reference distribution itself. We concluded that solving our novel formulation produces a solution that performs well on the reference distribution, is locally distributionally robust, and is quick in terms of computation time. For the Saturate with Preference setting, we showed that our algorithm consistently outperforms the standard Submodular Saturation Algorithm in terms of the performance on the objective functions with the highest assigned preference. For the application to the online submodular optimization setting, we demonstrated that our algorithm achieves a comparable utility to the regular method of solving each observed objective function at the moment of its observation, with a comparable wall-clock time taken. However, it does so by using a much smaller total number of distinct satellites. Finally, for a more general, real-life application of the proposed relative entropy-regularized algorithm, we tackled an image summarization task based on contemporary neural network usage.

### v3

@@ 6.2 Saturate with Preference for Satellite Selection
We now turn to the assessment of Algorithm 1, Saturate with Preference. The general setting of the experiments remains similar to the previous subsection, with the same constellation parameters of $75^{\circ}:240/12/1$, and the same objective functions $f_{t,1},\ldots,f_{t,6}$. We compare the performance of the selection made by Saturate with Preference, with that of the unmodified SSA, by looking at the values achieved by the two objective functions that are assigned the highest weight by the random sampling of $Q$. In essence, this allows us to evaluate whether adding the element of preference to SSA works as intended. Indeed, Figure 5 demonstrates that for the two objective functions with the highest priority, Saturate with Preference leads to the selection of a subset that consistently achieves a higher utility score in comparison to SSA. Indeed, Figure 5 demonstrates that Saturate with Preference consistently outperforms SSA in terms of the performance on the objective functions with the highest assigned priority.
Fig. 5: The average performances over fifteen runs of Saturate with Preference in comparison to SSA on the two objective functions with the highest assigned weight. The highlighted areas indicate one tenth standard deviation. The results have been put through a moving average filter with window size 66.
@@ 6.3 Application to Online Submodular Optimization
In this subsection, we assess the performance of (66), which we will call the time-robust (TR) formulation for the sake of brevity, as detailed in Section 5. In summary, the TR formulation aims, using a combination of a momentum-like weighing scheme of the time-varying objective functions in an online setting, along with the idea of relative-entropy regularization, to reuse the same selections made in one step over multiple time steps. In this way, it aims to be more cost-efficient in settings where the selection of more diverse elements incurs additional costs.
As a baseline, we choose the standard approach of treating each individual objective function $f_{t}$ observed at time step $t$ as a separate problem and solve it using the Stochastic Greedy algorithm, without any consideration for the conservation of solutions over time steps. Figure 6 demonstrates the comparison of our proposed approach with the time-window size $t_{w}=5$ against the standard approach, in terms of the utility achieved, the wall-clock time taken, and the total number of distinct elements used, i.e., $\lvert\bigcup_{t=0}^{T}\mathcal{S}_{t}\rvert$, where $\mathcal{S}_{t}$ represents the solution constructed at time step $t$ of the simulation. The results indicate that TR achieves a comparable utility in comparison to the standard approach. The wall-clock time taken fluctuates in the TR formulation, since with a time-window size of $5$, the algorithm only observes during four of every five time steps and plays its solution, performing all function evaluations on every fifth time step, although the total wall-clock time taken is in the same range as the standard approach. However, the total number of distinct elements chosen by the TR formulation is much lower, using less than half the number of distinct elements in comparison to the standard approach.
Fig. 6: The average performances over $15$ runs of the two algorithms in the online task detailed in Section 6.3. The comparison criteria are the utility achieved, the wall-clock time taken, and the total number of distinct elements used. Regular represents the standard approach of solving each objective function $f_{t}$ observed at time step $t$ independently of the previous objective functions observed, using the Stochastic Greedy algorithm. TR represents the time-robust formulation detailed in Section 5, with a time-window size $t_{w}=5$. The highlighted areas indicate one standard deviation. The results have been put through a moving average filter with size 66.
@@ 6.4 Practical Application: Image Summarization
As a final demonstration using a more up-to-date application involving machine learning and signal processing, we apply the proposed method to image summarization. We tackle the case detailed in [43], which, in short, involves selecting the most representative $K\leq\lvert\mathcal{N}\rvert$ images out of a dataset $\mathcal{N}=[\lvert\mathcal{N}\rvert$] of indexed images. More formally, for a selection $\mathcal{S}\subseteq\mathcal{N}$ of images, one can evaluate the utility $f_{i}(\mathcal{S})$ of each image $i\in\mathcal{N}$ through its similarity to the closest image in the selection by:
$f_{i}(\mathcal{S})=1-\min_{e\in\mathcal{S}}d(i,e),$ (69)
where $d(i,e)$ designates some distance between images $i$ and $e$. Replicating the setting in the aforementioned work [43], we use the $\lvert\mathcal{N}\rvert=819$-image Pokemon dataset[12], using image embeddings calculated with an AlexNet[55] trained on the ImageNet dataset[15]. We use the normalized cosine distance
$d(i,e)=\frac{v_{i}\cdot v_{e}}{\lVert v_{i}\rVert\lVert v_{e}\rVert}+\min_{(e,e^{{}^{\prime}})\in\mathcal{N}^{2}}\frac{v_{e}\cdot v_{e^{{}^{\prime}}}}{\lVert v_{e}\rVert\lVert v_{e^{{}^{\prime}}}\rVert}.$ (70)
The average results obtained from $15$ runs of the algorithms, where in each the reference distribution $Q$ is the uniform distribution, are demonstrated in Figure 7. We observe that our proposed Local algorithm outperforms SSA for nearly all values of the cardinality bound $K$ on the reference distribution and on the local worst-case distribution, while also taking significantly less computation time, as evidenced by the results of the previous experiments. Regarding the performance on the worst-case task, SSA only outperforms our algorithm within the cardinality constraint range of $K=1$ to $K=5$, but then both algorithms show virtually the same performance, again, with the Local algorithm being much less computationally expensive.
Fig. 7: The average performances over $15$ runs of the two algorithms focused on optimizing the global worst-case task, and the local worst-case distribution as guided by the reference distribution, evaluated on the three criteria of reference distribution performance, worst-case task performance, and local worst-case distribution performance in the image summarization task of Section 6.4. Local represents the relative-entropy-regularized Stochastic Greedy algorithm solving our novel formulation, aiming for local worst-case distributional robustness in the neighborhood of the reference distribution. Saturate (Global) represents SSA, aiming for global worst-case task robustness. The highlighted areas indicate one standard deviation. The results have been put through a moving average filter with window size 66.
@@ 7 Conclusion
We proposed the novel formulation of (14) to find a solution that is locally distributionally robust in the neighborhood of a reference distribution which assigns an importance score to each task in a multi-task objective, using various statistical distances as a regularizer. The $L^{1}$ and $L^{\infty}$ metrics led to the proposal of the Saturate with Preference algorithm, which incorporates an element of preference into the standard SSA. Afterward, we demonstrated that using relative entropy as a regularizer, through duality, we can show that this novel formulation is equivalent to (25). Then, we proved that this dual formulation gives rise to the composition of a monotone increasing function with a normalized, monotone nondecreasing submodular function, which can be optimized with standard methods such as the Stochastic Greedy algorithm, enjoying theoretical guarantees. We proposed an application of the relative-entropy-regularized objective to online submodular optimization, through the use of a momentum-like weighing of the objective functions observed over time steps. We then experimentally corroborated our results for all three of the proposed settings, motivated by a practical scenario involving a sensor selection problem within a simulation of LEO satellites, using weak submodular functions. For the relative-entropy regularization setting, we compared our algorithm with two other algorithms focused on optimizing the performance of the worst-case task, and on directly optimizing the performance reference distribution itself. We concluded that solving our novel formulation produces a solution that performs well on the reference distribution, is locally distributionally robust, and is cheap in terms of computation time. For the Saturate with Preference setting, we showed that our algorithm consistently outperforms the standard SSA in terms of the performance on the objective functions with the highest assigned preference. For the application to the online submodular optimization setting, we demonstrated that our algorithm achieves a comparable utility to the regular method of solving each observed objective function at the moment of its observation, with a comparable wall-clock time taken. However, it does so by using a much smaller total number of distinct satellites. Finally, for a more general, real-life application of the proposed relative-entropy-regularized algorithm, we tackled an image summarization task based on contemporary neural network usage.